39. Suppose a researcher wants to test the hypothesis that p = 0.37 versus the alternative that p ≠ 0.37 using data from a random sample of 29 people.  We calculate the standardized test statistic to be 1.84.  Which of the following best describes the p-value?

a.

0.025 < p-value < 0.05

b.

0.05 < p-value < 0.10

c.

p-value = 0.9671

d.

p value = 0.0658

e.

p-value = 0.0329
Answer #1

solution

Solution :

This is the two tailed test .

The null and alternative hypothesis is

H0 : p =0.37

Ha : p \neq 0.37

Test statistic = z =1.84

P(z >1.84 ) = 1 - P(z <1.84 ) = 1-0.9671=0.0329

P-value = 2*0.0329=0.0658